OpenStudy (anonymous):

Solve. -7/3n= -42 n=

OpenStudy (md152727):

n=18 Explanation:

OpenStudy (anonymous):

@pooja195 what do u think?

pooja195 (pooja195):

Multiply both sides by the reciprocal -7/3=3/-7 <----reciprocal -7/3 x 3/-7=n n=? 3/-7 x -42=?

OpenStudy (md152727):

\[3/-7*-7/3*n==3/-7*-42 so n=18\]

OpenStudy (anonymous):

thx last one guys